# A recursive approach for the determination of the nice sections of width three having a 4-crown stack as retract

### v1

@@ 3.2 Four-crown stacks as retracts of crowned sections
$D$ $\subset P(k+1\rightarrow h_{P}),$
@@
$\>\textrm{and}\>\quad\quad\quad\quad t$ $:P(0\rightarrow k)\rightarrow T\>\textrm{ is a retraction}\>$
$\>\textrm{with }\>S\>\textrm{ and }\>T$ being 2-antichains or 4-crown stacks,
@@
$\>\textrm{and}\>\quad\quad\quad\quad t$ $:P(k\rightarrow h_{P})\rightarrow T\>\textrm{ is a retraction}\>$
$\>\textrm{with }\>S\>\textrm{ and }\>T$ being 2-antichains or 4-crown stacks.
In order to avoid repetitions, we use the symbols $S$ and $T$ in what follows always as in this definition: for a retractive down-split $(k,D,s,t)$ of $P\in\mathfrak{C}$, we always have $S=s[P(k+1\rightarrow h_{P})\setminus D]$ and $T=t[P(0\rightarrow k)]$, and correspondingly for a retractive up-split. In a retractive down-split $(k,D,s,t)$, the poset $T$ is thus a retract of the lower segment of $P$ ending with level $k$ and $D$ is a down-set, whereas in a retractive up-split $(k,U,s,t)$, the poset $T$ is a retract of the upper segment of $P$ starting with level $k$ and $U$ is an up-set. The reader will observe that in a retractive down-split $(h_{P}-1,D,s,t)$ and a retractive up-split $(1,U,s,t)$ we must have $D\leq 1$ and $U\leq 1$, respectively, because $S$ contains at least two points.
Theorem 3.4.
A poset $P\in\mathfrak{C}$ has a 4-crown stack as retract iff there exists a retraction $r:P\rightarrow R$ of $P$ onto a 4-crown stack $R$ for which
  • a retractive down-split $(k,D,s,t)$ exists with
    $\begin{split}R&=T\oplus S,\\ r|_{P(k+1\rightarrow h_{P})\setminus D}^{P(k+1\rightarrow h_{P})\setminus D}&=s,\\ r|_{P(0\rightarrow k)}^{P(0\rightarrow k)}&=t,\\ r[D]&\subseteq T,\end{split}$ (1)
  • or a retractive up-split $(k,U,s,t)$ exists with
    $\begin{split}R&=S\oplus T,\\ r|_{P(0\rightarrow k-1)\setminus U}^{P(0\rightarrow k-1)\setminus U}&=s,\\ r|_{P(k\rightarrow h_{P})}^{P(k\rightarrow h_{P})}&=t,\\ r[U]&\subseteq T.\end{split}$ (2)
The retraction of the poset on the left in Figure 2 yields for $k=0$ a retractive down-split fulfilling (1) and for $k=3$ a retractive up-split fulfilling (2). For the retraction of the poset on the right of the figure, the suitable level indices are $0,2$ and $1,3$, respectively. A retractive split of a crowned section $P$ and a retraction of $P$ onto a 4-crown stack are called matching if they are coupled by (1) or (2).
Only direction $\Rightarrow$” of the theorem has to be proven. We agree on that for the rest of this section the poset $P$ is always a crowned section and $r:P\rightarrow R$ is a retraction onto a 4-crown stack. We start the proof with two simple observations. Firstly, the down-sets $r^{-1}(a)$, $a\in R(0)$, are pairwise disjoint, hence
$R(0)\cap P(0)\not$$\emptyset$. (3)
Secondly,
@@
Let $r[P(0\rightarrow k)]=R(0\rightarrow\ell)\subset P(0\rightarrow k)$ with $\ell<h_{R}$. A matching down-split $(k,D,s,t)$ is given by $t:=r|_{P(0\rightarrow k)}^{P(0\rightarrow k)}$,
$D:=\left\{x\in P(k+1\rightarrow h_{P})\>\left|\>r(x)\in R(0\rightarrow\ell)\right.\right\},\quad s:=r|_{P(k+1\rightarrow h_{P})\setminus D}^{P(k+1\rightarrow h_{P})\setminus D}.$
In the next two propositions we see that two types of level sets of $R$ always give raise to a matching retractive split of $P$: the level sets of $R$ contained in single level sets of $P$, and the level sets of $R$ with points in “maximal possible distance” in $P$. For examples, see Figure 2.
Corollary 3.6.
There exists a matching retractive split of $P$ if there exists a level set of $R$ contained in a single level set of $P$.
Proof.
Let $\ell$ and $k$ be level indices of $R$ and $P$, respectively, with $R(\ell)\subset P(k)$ and let $z\in P(k)\setminus R(\ell)$. Three cases are possible:
  1. 1.
    $\ell\in[1,h_{R}-1]$: Then $k\in[1,h_{P}-1]$. Depending on $\lambda(r(z))\leq k$ or $\lambda(r(z))\geq k$, Lemma 3.5 and its dual yield a matching retractive split.
  2. 2.
    $\ell=0$: Then $k=0$ due to (3). If $r(z)\notin R(0)$, define
    $t:=r|_{P(1\rightarrow h_{P})}^{P(1\rightarrow h_{P})},\quad U:=\{z\},\quad s:=r|_{P(0)\setminus U}^{P(0)\setminus U}.$
    For $x\in P(1\rightarrow h_{P})$, we have $R(0)<x$ or $z<x$, hence $r(x)\in P(1\rightarrow h_{P})$. Now apply the dual of Lemma 3.5. And in the case of $r(z)\in R(0)$, apply the lemma on
    $t:=r|_{P(0)}^{P(0)},\quad D:=\left\{x\in P(1\rightarrow h_{P})\>\left|\>r(x)\in P(0)\right.\right\},\quad s:=r|_{P(1\rightarrow h_{P})\setminus D}^{P(1\rightarrow h_{P})\setminus D}.$
  3. 3.
    $\ell=h_{R}$: Dual to the previous case.
Lemma 3.7.
Let $R(\ell)=\{a,b\}$ with $\lambda(a)<\lambda(b)$. There exists a matching retractive split of $P$ if
  1. 1.
    $\ell=0$ and $P(0,\lambda(b)+1)\simeq 33$;
  2. 2.
    $\ell=h_{R}$ and $P(\lambda(a)-1,h_{P})\simeq 33$;
  3. 3.
    $0<\ell<h_{R}$ and $P(\lambda(a)-1,\lambda(b))\simeq 33\simeq P(\lambda(a),\lambda(b)+1)$.
Proof.
1) We have $\lambda(a)=0$ according to (3). Additionally, there exists a point $x\in P(0)$ with $r(x)=b$, hence $R(0)<r(y)$ for all $y\in P(\lambda(b)+1\rightarrow h_{P})$. Now apply the dual of Lemma 3.5.
2) Dual to the previous proposition.
3) Let $k:=\lambda(a)$, $j:=\lambda(b)$, and $x\in P(k)\setminus\{a\}$. If $r(x)\in\{b\}\cup R(\ell+1\rightarrow h_{R})$, then $P(k,j+1)\simeq 33$ yields $R(\ell)<r(z)$ for all $z\in P(j+1\rightarrow h_{P})$, and the dual of Lemma 3.5 yields a matching up-split $(j+1,U,s,t)$.
Now assume $r(x)\in R(0\rightarrow\ell-1)\cup\{a\}$ for all $x\in P(k)\setminus\{a\}$. Then $r(z)\in R(0\rightarrow\ell-1)\cup\{a\}$ for all $z\in P(0\rightarrow k-1)$, and $P(k-1,j)\simeq 33$ additionally yields $r(z)<b$, hence $r(z)\in P(0\rightarrow k-1)$. According to Lemma 3.5, there exists a matching down-split (k1,D,s,t)(k-1,D,s,t).
By means of Corollary 3.6, we easily prove direction $\Rightarrow$” of Theorem 3.4. If $r^{\prime}:P\rightarrow R^{\prime}$ is a retraction of $P$ onto a 4-crown stack, then Lemma 2.1 ensures the existence of a retraction $r:P\rightarrow R$ with $R\simeq R^{\prime}$ being a spanning retract, hence $R(0)\subset P(0)$ and $R(h_{R})\subset P(h_{P})$. Now Corollary 3.6 yields the existence of a retractive split of $P$ coupled with $r$ by (1) or (2), respectively.
3.3 A recursive approach
Let $\mathfrak{S}$ be a subclass of $\mathfrak{C}$ closed under duality, e.g., the subclass of nice sections of width three. Theorem 3.4 states that $P\in\mathfrak{S}$ having a 4-crown stack as retract requires a proper lower segment $P(0\rightarrow k)$ or a proper upper segment $P(k\rightarrow h_{P})$ having a 2-antichain or a 4-crown as retract. This gives raise to a recursive approach for the determination of the sections in $\mathfrak{S}$ having a 4-crown stack as retract.
For a retractive down-split or up-split, we need a $t$-base $V$ and an $s$-base $W$ of a poset $P\in\mathfrak{S}$ of width 3, i.e.:
@@ 3.3 A recursive approach
Assume that we have listed all lower segments up to height $n$ of sections contained in $\mathfrak{S}$ which have a 2-antichain or a 4-crown stack as retract. Given a poset $P\in\mathfrak{S}$ of height $n+1$, we can quickly identify all candidates for $t$-bases in $P$ by checking which of the lower segments in our list are lower segments $P(0\rightarrow k)$ of $P$ and which of their duals are upper segments $P(k\rightarrow n+1)$ of $P$. Of course, $P(0)$ and $P(n+1)$ are always candidates for a $t$-base. According to Theorem 3.4, the poset $P$ has a 4-crown stack as retract iff one of the corresponding segments $P(k+1\rightarrow n+1)$ or $P(0\rightarrow k-1)$ provides an $s$-base fulfilling (1) or (2), respectively.
@@ 4.1 Definition and isomorphism types
For an integer $\eta\in\mathbb{N}$, the symbols $\mathfrak{C}_{\eta}$ and $\mathfrak{N}_{\eta}$ denote the classes of crowned sections and nice sections with horizon $\eta$ and height $\geq 2$. The set $\mathfrak{C}_{1}$ is empty, and the only crowned sections not contained in $\mathfrak{C}_{\eta}$ for any $\eta\geq 2$ are the 6-crowns. For the rest of this article we focus on $\eta=2$, and for this horizon, we can skip the discrimination between nice sections and crowned sections:
@@
Let $P\in\mathfrak{C}_{2}$. We have to prove that $P$ does not contain an irreducible point. Let $x\in P(k)$, $k\in[0,h_{P}-1]$. If $P(k,k+1)$ is a 6-crown, $x$ has two upper covers. Let thus $P(k,k+1)\simeq 3C$, hence $k\leq h_{P}-2$ and $x<P(k+2)$. The layer $P(k+1)$ contains an upper cover of $x$, and depending on the type of $P(k+1,k+2)$, the layer $P(k+2)$ contains one or even two additional upper covers of $x$. The argument for lower covers is dual.
@@ 4.2 Properties of retracts
We start our investigation with three simple observations; in particular implication (4) will play an important role in several proofs because it yields via Lemma 3.5 a matching retractive up-split:
@@
2) Let $\ell\in[0,h_{R}]$ with $R(\ell)=2$ and $0<\rho:=\max\left\{\lambda(x)\>\left|\>x\in R(\ell)\right.\right\}<h_{P}$. Then
$\begin{split}&\forall\;a\in R(\ell):\;r^{-1}(a)\cap P(\rho-1)\not=\emptyset\\ \>\textrm{implies}\>\quad&r[P(\rho+1\rightarrow h_{P})]=R(\ell+1\rightarrow h_{R})\subseteq P(\rho+1\rightarrow h_{P}).\end{split}$ (4)
@@
The first proposition follows from the definition of the horizon and (3). For the second one, observe that the assumption implies $a\leq r(y)$ for all $a\in R(\ell)$ and all $y\in P(\rho+1\rightarrow h_{P})$, hence $R(\ell)<r(y)$.
@@
$\Rightarrow$”: Let $R(0)=\{a,b\}$ with $\lambda(a)\leq\lambda(b)$. Due to the first part of Corollary 4.3, we have $R(0)\subset P(0,1)$. In the case of $\lambda(b)=1$, the equation $r[P(2\rightarrow h_{P})]=R(1\rightarrow h_{P})$ follows with implication (4). In the case of $\lambda(b)=0$, the relation $R(0)<P(2\rightarrow h_{P})$ enforces $r[P(2\rightarrow h_{P})]\subseteq R(1\rightarrow h_{P})$, and $P(0,1)\simeq 3C$ yields $R(1)\subset P(2\rightarrow h_{P})$.
@@ 4.3 Retractive up- and down-splits
Let $P\in\mathfrak{N}_{2}$ and let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$. Due to the horizon of $P$, every level set of $R$ is either contained in a single level set of $P$ or consists of points belonging to consecutive level sets of $P$. Now Corollary 3.6 and Lemma 3.7 tell us that every level-set of $R$ gives raise to matching retractive split of $P$. On the other hand, for a given retractive split of $P$ it is easy to check if there exists a matching retraction of $P$ onto a 4-crown stack:
@@
Let $P\in\mathfrak{N}_{2}$. For a retractive down-split $(k,D,s,t)$ of $P$, the existence of a retraction $r:P\rightarrow R$ onto a 4-crown stack $R$ coupled with $(k,D,s,t)$ by (1) is equivalent to
@@
which implies $D\subseteq P(k+1)$, and for a retractive up-split $(k,U,s,t)$ of $P$, there exists a retraction $r:P\rightarrow R$ onto a 4-crown stack $R$ coupled with $(k,U,s,t)$ by (2) iff
@@
Let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$ coupled with a retractive up-split $(k,U,s,t)$ by (2). Due to $r[P(k\rightarrow h_{P})]=T$, the second part of Condition (6) is necessary for $r[U]\subseteq T$. The addendum follows because of $P(0\rightarrow k-2)<_{P}P(k\rightarrow h_{P})\supset T(0)$.
Now assume that (6) holds for $(k,U,s,t)$ and let $u\in U\subseteq P(k-1)$. There exists a point $\tau(u)\in T(0)$ with $u\not<_{P}p$ for all $p\in t^{-1}(\tau(u))$. We define $\rho(u)$ as the single point contained in $T(0)\setminus\{\tau(u)\}$. Now we define a mapping $r:P\rightarrow P$ by setting for $x\in P$
@@
Let $S(h_{S})=\{a,b\}$ and $T(0)=\{v,w\}$. The assumptions imply $a\in P(k-1)$ and $v\in P(k+1)$. Assume $P(k,k+1)\simeq 3C$. According to Lemma 2.1, we can assume $S(h_{S})\subset P(k-1)$ without loss of generality. Let $P(k-1)=\{a,b,z\}$.
@@
No point of $P(k)\setminus U$ is above $a$ and the point $b$ is the only maximal point of $S\setminus\{a\}$. Due to $s^{-1}(a)=\{a\}$, the mapping
$s^{\prime}:P(0\rightarrow k)\setminus U$ $\rightarrow S,$
$x$ $\mapsto\begin{cases}s(x),&\>\textrm{if }\>x\in P(0\rightarrow k-1),\\ b,&\>\textrm{if }\>x\in P(k)\setminus U\end{cases}$
@@ 5 Application
With the recursive approach described in Section 3.3, we will now identify all posets $P\in\mathfrak{N}^{LS}_{2}$ with height up to six having a 4-crown stack as retract. For some of them, e.g. the 6-crown stacks, we already know the result. We include them in our investigation because our purpose is to test our approach with as many nice sections as possible.
We encode the lower segments not being 3-antichains by the sequence of their level-pair types (cf. Theorem 4.2). 1 indicates a 6-crown, 0 a type $3C$. The lower segment 111 is thus the 6-crown stack of height 3, whereas 110 indicates the lower segment starting with two 6-crowns followed by a single $3C$.
@@
Table 1: The lower segments from $\mathfrak{N}^{LS}_{2}$ with height from one to six. The letters “y” and “n” indicate if a segment has 4-crown stack as retract or not. Additional explanation in text.
@@
For the posets with a final 0, we can decide by applying the dual of Corollary 4.4 referring to previous results in Table 1. As an example, the posets $110$ and $100$ do not have a 4-crown stack as retract, because the 6-crown $P(0,1)$ does not have a suitable retract. And $1010$ and $1000$ have a 4-crown stack as retract because 10 is marked with “y”.
@@
Figure 5: a) The lower segment $10$ and a retraction $r$ onto a 4-crown with $r^{-1}(v)=\{v\}$ for a point $v$ from the top level of the 4-crown. b) Illustrations for the case discrimination in the proof of Criterion 5.1.5. The hollow dots belong to $S(0)$.
@@
Let $P\in\mathfrak{N}_{2}$ with $h_{P}\geq 3$.
@@
    For every retractive down-split $(h_{P}-1,D,s,t)$ we have $t[P(0\rightarrow h_{P}-3)]=T(0\rightarrow h_{T}-1)$. In particular, the choice of $P(h_{P})$ as an $s$-base implies that the poset $P(0\rightarrow h_{P}-3)$ has a 2-antichain or a 4-crown tower as retract.
@@
    For the choice of $P(h_{P}-1,h_{P})$ as $s$-base, the set $D\subseteq P(h_{P}-1)$ must contain at least two points.
@@
    The level-pair $P(h_{P}-1,h_{P})$ is not a candidate for an $s$-base if the level-pair $P(h_{P}-2,h_{P}-1)$ is a 6-crown.
@@
    $P(1\rightarrow h_{P})$ can serve as an $s$-base only together with a subset $D\subset P(1)$ containing at most a single point.
@@
    If $P(1,2)$ is a 6-crown, then $P(1\rightarrow h_{P})$ being an $s$-base implies that $P(3\rightarrow h_{P})$ has a 2-antichain or a 4-crown stack as retract.
@@
1. The first condition in (5) yields $T(h_{T})\not\subset P(h_{P}-1)$. Now the dual of Implication (4) yields $t[P(0\rightarrow h_{P}-3)]=T(0\rightarrow h_{T}-1)$.
@@
3. Is implied by the second criterion, because if $P(h_{P}-2,h_{P}-1)$ is a 6-crown, the second condition in (5) can be fulfilled for at most a single point $d\in P(h_{P}-1)$.
4. Trivial.
5. We discriminate three cases (cf. Figure 5b). If $S(0)$ contains a point $a\in P(1)$, then it has to contain a point $b\in P(2)$, too, because of the first condition in (5). At least one of the points in $P(1)\setminus\{a\}$ has to be mapped to $b$ (fourth criterion), thus $s[P(3\rightarrow h_{P})]=S(1\rightarrow h_{S})$ due to implication (4).
Now assume that $S(0)$ does not contain a point of $P(1)$. If $S(0)$ contains a point $v\in P(3\rightarrow h_{P})$, then at least two points in $P(1)$ are below $S(0)$ in contradiction to the fourth criterion. And in the case of $S(0)\subset P(2)$, exactly one point $x\in P(1)$ is below $S(0)$ and has to be sent to $P(0)$. From the two remaining points in $P(1)$ each is under a single different point of $S(0)$. Due to the fourth criterion, they have to be mapped onto $S(0)$, and $s[P(3\rightarrow h_{P})]=S(1\rightarrow h_{S})$ follows again with implication (4).
The rightmost poset in Figure 7 confirms that we can neither drop the condition $P(h_{P}-2,h_{P}-1)$ is a 6-crown” in Criterion 5.1.3 nor the condition “$P(1,2)$ is a 6-crown” in Criterion 5.1.5.
@@
Now assume $D=\{d,y\}$ with $d\in P(0)$ and $d\lessdot y$. The poset $P\setminus D$ is shown in Figure 6c. If a retraction as described in the lemma exists, it cannot be extended to $P\setminus\{d\}$. The upper covers $a$ and $b$ of $y$ are thus mapped to different points of $R(0)$. But that is impossible because of $P(0)\setminus\{d\}<\{a,b\}$.

### v2

@@ 3.3 A recursive approach
$D$ $\subset P(k+1\rightarrow h_{P})\>\textrm{ is a down-set,}\>$
@@
$\>\textrm{and}\>\quad\quad\quad\quad t$ $:P(0\rightarrow k)\rightarrow T\>\textrm{ is a retraction,}\>$
$\>\textrm{where }\>S\>\textrm{ and }\>T$ are 2-antichains or 4-crown stacks,
@@
$\>\textrm{and}\>\quad\quad\quad\quad t$ $:P(k\rightarrow h_{P})\rightarrow T\>\textrm{ is a retraction,}\>$
$\>\textrm{where }\>S\>\textrm{ and }\>T$ are 2-antichains or 4-crown stacks.
In order to avoid repetitions, we use the symbols $S$ and $T$ in what follows always as in this definition: for a retractive down-split $(k,D,s,t)$ of $P\in\mathfrak{C}$, we always have $S=s[P(k+1\rightarrow h_{P})\setminus D]$ and $T=t[P(0\rightarrow k)]$, and correspondingly for a retractive up-split. In a retractive down-split $(k,D,s,t)$, the poset $T$ is thus a retract of the lower segment of $P$ ending with level $k$ and $D$ is a down-set in $P(k+1\rightarrow h_{P})$, whereas in a retractive up-split $(k,U,s,t)$, the poset $T$ is a retract of the upper segment of $P$ starting with level $k$ and $U$ is an up-set in $P(0\rightarrow k-1)$. The reader will observe that in a retractive down-split $(h_{P}-1,D,s,t)$ and a retractive up-split $(1,U,s,t)$ we must have $D\leq 1$ and $U\leq 1$, respectively, because $S$ contains at least two points.
We have seen in Section 3.2 that a crowned section has a 4-crown stack as retract iff it has such a retract with a split level. In terms of retractive splits this means
@@
There exists a retraction of $P\in\mathfrak{C}_{2}$ onto a 4-crown stack iff
  • a retractive down-split $(k,D,s,t)$ of $P$ exists for which $T(h_{T})<S(0)$ and for which the mapping
    $p$ $\mapsto\begin{cases}s(p),&\>\textrm{if }\>p\in P(k+1\rightarrow h_{P})\setminus D,\\ t(p),&\>\textrm{if }\>p\in P(0\rightarrow k)\end{cases}$
    can be extended to a retraction $r:P\rightarrow T\oplus S$,
  • or a retractive up-split $(k,U,s,t)$ of $P$ exists for which $S(h_{S})<T(0)$ and for which the mapping
    $p$ $\mapsto\begin{cases}s(p),&\>\textrm{if }\>p\in P(0\rightarrow k-1)\setminus U,\\ t(p),&\>\textrm{if }\>p\in P(k\rightarrow h_{P})\end{cases}$
    can be extended to a retraction $r:P\rightarrow S\oplus T$.
We call a retraction $r$ and a retractive split matching if they are coupled in this way.
This result gives raise to a recursive approach to determine the crowned sections having a 4-crown stack as retract. Let $\mathfrak{S}$ be a subclass of $\mathfrak{C}$ closed under duality, e.g., the subclass of nice sections of width three. For a retractive split of a poset $P\in\mathfrak{S}$, we need a $t$-base $V$ and an $s$-base $W$ of a poset $P\in\mathfrak{S}$ of width 3, i.e.,
@@
Assume that we have listed all lower segments up to height $n$ of sections contained in $\mathfrak{S}$ which have a 2-antichain or a 4-crown stack as retract. Given a poset $P\in\mathfrak{S}$ of height $n+1$, we can quickly identify all candidates for $t$-bases in $P$ by checking which of the lower segments in our list are lower segments $P(0\rightarrow k)$ of $P$ and which of their duals are upper segments $P(k\rightarrow n+1)$ of $P$. Of course, $P(0)$ and $P(n+1)$ are always candidates. The poset $P$ has a 4-crown stack as retract iff one of the corresponding segments $P(k+1\rightarrow n+1)$ or $P(0\rightarrow k-1)$ provides an $s$-base which can be combined with the $t$-base as described in Lemma 3.11.
@@ 4.1 Definition and isomorphism types
For an integer $\eta\in\mathbb{N}$, the symbols $\mathfrak{C}_{\eta}$ and $\mathfrak{N}_{\eta}$ denote the classes of crowned sections and nice sections with horizon $\eta$ and height at least 2. The set $\mathfrak{C}_{1}$ is empty, and the only crowned sections not contained in $\mathfrak{C}_{\eta}$ for any $\eta\geq 2$ are the 6-crowns. For the rest of this article we focus on $\eta=2$, and for this horizon, we can skip the discrimination between nice sections and crowned sections:
@@
Let $P\in\mathfrak{C}_{2}$. We have to prove that $P$ does not contain an irreducible point. Let $x\in P(k)$, $k\in[0,h_{P}-1]$. For the upper covers of $x$, only the case $P(k,k+1)\simeq 3C$ is of interest which yields $k\leq h_{P}-2$. One upper cover of $x$ is contained in $P(k+1)$, and depending on the type of $P(k+1,k+2)$, the layer $P(k+2)>x$ contains one or even two additional upper covers of $x$. The argument for lower covers is dual.
@@ 4.2 Properties of retracts
We start our investigation with three simple observations:
@@
2) Let $\ell\in[0,h_{R}]$ with $R(\ell)=2$ and $0<\rho:=\max\lambda[R(\ell)]<h_{P}$. We have $r[P(\rho+1\rightarrow h_{P})]=R(\ell+1\rightarrow h_{R})$ if
$r^{-1}(a)\cap P(\rho-1)\not=\emptyset\quad\>\textrm{for all }\>a\in R(\ell).$ (2)
In particular, the level set $R(\ell+1)$ is an up-split level.
@@
The first proposition follows from the definition of the horizon and (1). For the second one, observe that the assumption implies $a\leq r(y)$ for all $a\in R(\ell)$ and all $y\in P(\rho+1)$, hence $R(\ell)<r(y)$.
@@
$\Rightarrow$”: Let $R(0)=\{a,b\}$ with $\lambda(a)\leq\lambda(b)$. Due to the first part of Corollary 4.3, we have $R(0)\subset P(0,1)$. In the case of $\lambda(b)=1$, the equation $r[P(2\rightarrow h_{P})]=R(1\rightarrow h_{P})$ follows with (2). In the case of $\lambda(b)=0$, the relation $R(0)<P(2\rightarrow h_{P})$ enforces $r[P(2\rightarrow h_{P})]\subseteq R(1\rightarrow h_{P})$, and $P(0,1)\simeq 3C$ yields $R(1)\subset P(2\rightarrow h_{P})$.
@@ 4.3 Retractive up- and down-splits
Let $P\in\mathfrak{N}_{2}$ and let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$. Due to the horizon of $P$, every level set of $R$ is either contained in a single level set of $P$ or consists of points belonging to consecutive level sets of $P$. Now Lemma 3.4 and Lemma 3.6 tell us that every level-set of $R$ gives raise to matching retractive split of $P$. On the other hand, for a given retractive split of $P$ it is easy to check if there exists a matching retraction of $P$ onto a 4-crown stack:
@@
Let $P\in\mathfrak{N}_{2}$. For a retractive down-split $(k,D,s,t)$ of $P$, the existence of a matching retraction $r:P\rightarrow T\oplus S$ is equivalent to
@@
which implies $D\subseteq P(k+1)$, and for a retractive up-split $(k,U,s,t)$ of $P$, there exists a matching retraction $r:P\rightarrow S\oplus T$ iff
@@
Let $r:P\rightarrow R$ be a retraction onto a 4-crown stack $R$ matching a retractive up-split $(k,U,s,t)$. Due to $r[P(k\rightarrow h_{P})]=T$, the second part of Condition (4) is necessary for $r[U]\subseteq T$. The addendum follows because of $P(k-2)<_{P}P(k)\supset T(0)$.
Now assume that (4) holds for $(k,U,s,t)$ and let $u\in U\subseteq P(k-1)$. There exists a point $\tau(u)\in T(0)$ with $u\not<_{P}p$ for all $p\in t^{-1}(\tau(u))$. We define $\rho(u)$ as the single point contained in $T(0)\setminus\{\tau(u)\}$. We define a mapping $r:P\rightarrow P$ by setting for $x\in P$
@@
Let $S(h_{S})=\{a,b\}$ and $T(0)=\{v,w\}$. The assumptions imply $a\in P(k-1)$ and $v\in P(k+1)$. Assume $P(k,k+1)\simeq 3C$ (the other case is dual). According to Lemma 2.1, we can assume $S(h_{S})\subset P(k-1)$ without loss of generality. Let $P(k-1)=\{a,b,z\}$.
@@
No point of $P(k)\setminus U$ is above $a$ and the point $b$ is the only maximal point of $S\setminus\{a\}$. Due to $s^{-1}(a)=\{a\}$, the mapping $s^{\prime}:P(0\rightarrow k)\setminus U\rightarrow S$ defined by
$s^{\prime}(x)$ $:=\begin{cases}s(x),&\>\textrm{if }\>x\in P(0\rightarrow k-1),\\ b,&\>\textrm{if }\>x\in P(k)\setminus U\end{cases}$
@@ 5 Application
With the recursive approach described in Section 3.3, we will now identify all posets $P\in\mathfrak{N}^{LS}_{2}$ with height up to six which have a 4-crown stack as retract. For some of them, e.g. the 6-crown stacks, we already know the result. We include them in the investigation because our purpose is to test our approach with as many nice sections as possible.
We encode the lower segments not being 3-antichains by the sequence of their level-pair types (cf. Theorem 4.2). The letter “1” indicates a 6-crown, “0” a type $3C$. The lower segment 111 is thus the 6-crown stack of height 3, whereas 110 indicates the lower segment starting with two 6-crowns followed by a single $3C$.
@@
Table 1: The lower segments in $\mathfrak{N}^{LS}_{2}$ with height from one to six. Explanation in text.
@@
For the posets with a final 0, we can apply the dual of Corollary 4.4 on previous results in Table 1. As an example, the posets $110$ and $100$ do not have a 4-crown stack as retract, because the 6-crown $P(0,1)$ does not have a suitable retract. And $1010$ and $1000$ have a 4-crown stack as retract because 10 is marked with “y”.
@@
Figure 7: a) The lower segment $10$ and a retraction $r$ onto a 4-crown with $r^{-1}(v)=\{v\}$ for a point $v$ from the top level of the 4-crown. b) Illustrations for the case discrimination in the proof of Criterion 5.1.2. The hollow dots belong to $S(0)$.
@@
Let $P\in\mathfrak{N}_{2}$ with $h_{P}\geq 3$. Assume that there exists a matching retraction $r$ for the retractive down-split $(k,D,s,t)$.
@@
    $k=0$ implies that $D\subset P(1)$ contains at most a single point.
@@
    $k=0$ and $P(1,2)$ being a 6-crown implies that the poset $P(3\rightarrow h_{P})$ has a 2-antichain or a 4-crown stack as retract.
@@
    $k=h_{P}-2$ implies that $D\subseteq P(k-1)$ contains at least two points.
@@
    $k=h_{P}-2$ and $P(h_{P}-2,h_{P}-1)$ being a 6-crown implies that the poset $P(0\rightarrow h_{P}-3)$ has a 2-antichain or a 4-crown stack as retract.
@@
    $k=h_{P}-1$ implies that the poset $P(0\rightarrow h_{P}-3)$ has a 2-antichain or a 4-crown stack as retract.
@@
1. Trivial.
2. We discriminate three cases (cf. Figure 7b). If $S(0)$ contains a point $a\in P(1)$, then it has to contain a point $b\in P(2)$, too, because of the first condition in (3). At least one of the points in $P(1)\setminus\{a\}$ has to be mapped to $b$ (first criterion), thus $s[P(3\rightarrow h_{P})]=S(1\rightarrow h_{S})$ due to (2).
Now assume that $S(0)$ does not contain a point of $P(1)$. If $S(0)$ contains a point $v\in P(3\rightarrow h_{P})$, then at least two points in $P(1)$ are below $S(0)$ in contradiction to the first criterion. And in the case of $S(0)\subset P(2)$, exactly one point $x\in P(1)$ is below $S(0)$ and has to be sent to $P(0)$. From the two remaining points in $P(1)$ each is under a single different point of $S(0)$. Due to the first criterion, they have to be mapped onto $S(0)$, and $s[P(3\rightarrow h_{P})]=S(1\rightarrow h_{S})$ follows again with (2).
@@
4. According to the third criterion, the set $D\subset P(h_{P}-1)$ contains at least two points. The second condition in (3) says that every point in $D$ is comparable with a single point of $T(h_{T})$ only which implies $T(h_{T})\subset P(h_{P}-2)$. A matching retraction maps $D$ onto $T(h_{T})$, and the dual of (2) yields the result.
5. The first condition in (3) yields $T(h_{T})\not\subset P(h_{P}-1)$. Now the dual of (2) yields $t[P(0\rightarrow h_{P}-3)]=T(0\rightarrow h_{T}-1)$.
The rightmost poset in Figure 9 confirms that we can neither drop the condition “$P(1,2)$ is a 6-crown” in Criterion 5.1.2 nor the condition $P(h_{P}-2,h_{P}-1)$ is a 6-crown” in Criterion 5.1.4.
@@
Now assume $D=\{d,y\}$ with $d\in P(0)$ and $d\lessdot y$. The poset $P\setminus D$ is shown in Figure 8c. If a retraction as described in the lemma exists, it cannot be extended to $P\setminus\{d\}$. The upper covers $a$ and $b$ of $y$ are thus mapped to different points of $R(0)$. But this is impossible because of $P(0)\setminus\{d\}<\{a,b\}$.